### v1

# Online Learning of Whittle Indices for Restless Bandits with Non-Stationary Transition Kernels

Abstract
We consider optimal resource allocation for restless multi-armed bandits (RMABs) in unknown, non-stationary settings. RMABs are PSPACE-hard to solve optimally, even when all parameters are known. The Whittle index policy is known to achieve asymptotic optimality for a large class of such problems, while remaining computationally efficient. In many practical settings, however, the transition kernels required to compute the Whittle index are unknown and non-stationary. In this work, we propose an online learning algorithm for Whittle indices in this setting. Our algorithm first predicts current transition kernels by solving a linear optimization problem based on upper confidence bounds and empirical transition probabilities calculated from data over a sliding window. Then, it computes the Whittle index associated with the predicted transition kernels. We design these sliding windows and upper confidence bounds to guarantee sub-linear dynamic regret on the number of episodes $T$, under the condition that transition kernels change slowly over time (rate upper bounded by $\epsilon=1/T^{k}$ with $k>0$). Furthermore, our proposed algorithm and regret analysis are designed to exploit prior domain knowledge and structural information of the RMABs to accelerate the learning process. Numerical results validate that our algorithm achieves superior performance in terms of lowest cumulative regret relative to baselines in non-stationary environments.
1 Introduction
Many sequential decision-making problems can be modeled as restless multi-armed bandits (RMABs). A decision maker needs to choose $M$ out of $N$ arms to activate at each time-slot. Each arm is modeled as a Markov decision process, and evolves stochastically according to two different transition kernels, depending on whether the arm is activated or not. At the beginning of each time-slot, the decision maker picks a subset of arms to be activated. The activated arms evolve according to their active Markov transition kernels, while the rest of the arms evolve according to their passive Markov transition kernels. At the end of the time-slot, the decision maker receives rewards from each arm, where rewards are functions of the current state and the action.
RMABs have a long history in resource allocation and operations research literature, starting with Whittle’s seminal work [1] in the 1980s. Over the past four decades, RMABs have been used to model and optimize resource allocation problems in a wide variety of domains such as wireless scheduling [2, 3, 4, 5, 6], machine monitoring and control [7, 8, 9], server scheduling [10], recommendation systems [11, 12], and health care [13, 14, 15, 16, 17]. In all of these applications, transition kernels can be unknown and non-stationary, i.e., the laws governing the evolution of states can drift over time. For example, consider a load balancing problem, where jobs arrive into a datacenter and a decision-maker assigns jobs to servers via a load balancer. The time required to finish a job at any server depends on its current load and how the load evolves over time. This evolution is typically random and time-varying since there are multiple load balancers and job streams contributing to the load at any given server within a large datacenter. Deciding which server to pick can then be formulated as an RMAB, but with non-stationary transition kernels.
Even when the transition kernels are known, developing an optimal policy for RMABs is PSPACE-hard [18] because of the combinatorial action space. The Whittle index policy [1] achieves asymptotic optimality [19, 20] for a special class of RMABs that satisfy the indexability property. Most prior works in in this space focus on stationary transition kernels (either known [21, 4, 5, 22, 12] or unknown [23, 24]), leaving a large gap for studying RMABs with unknown and non-stationary transition kernels. In addition, it is common in many applications to have prior knowledge regarding the transition kernels for some parts of the state space. For example, consider a wireless scheduling problem which aims to maximize information freshness in selecting which users (arms) to schedule. In this case, the state can be modeled using Age of Information (AoI) [25, 26] – a widely used metric for quantifying information freshness. Then, the AoI of an arm increases by one if the arm is not scheduled for transmission. Conversely, if the arm is scheduled, its AoI resets to one with the success probability of the transmission. Thus, the AoI will never increase by 2 or decrease to a value other than 1. Applying techniques from online and reinforcement learning naively would require attempting to learn the entire transition matrix, even though parts of it are known, leading to poor performance and inefficient regret guarantees.
In this paper, we pose the following research question: Can we develop a Whittle index-based online algorithm for RMABs with non-stationary transition kernels? To address this, we take prior domain knowledge for RMABs into account and provide online algorithms and regret guarantees that scale with the amount of uncertainty about the bandits.
Contributions: The main contributions of our paper can be summarized as follows:
  • We model non-stationarity of RMABs by assuming an arbitrary upper-bound $\epsilon$ on the total variational divergence between transition kernels in two consecutive episodes. Based on this, we propose an online learning algorithm which predicts the Whittle index by using a sliding window and upper confidence bound approaches. The sliding window and the upper confidence bound are determined as functions of the parameter $\epsilon$. Moreover, our solution incorporates prior domain knowledge and structural information of the RMABs. In particular, our algorithm employs different window sizes and upper confidence bounds for the subset of state-action pairs with unknown yet stationary transition kernels compared to state-action pairs with non-stationary transition kernels.
  • We rigorously characterize an upper bound on the dynamic regret of our algorithm. Our analysis demonstrates the impact of the parameter $\epsilon$ and the window size $w$ on the dynamic regret, allowing design of $w$ as a function of $\epsilon$. Given $\epsilon\leq 1/T^{k}$, our design ensures sub-linear dynamic regret $O(\sqrt{\mathrm{log}(T)}T^{r})$ on the total number of episodes $T$ with $k>0$, $r=\max\{(3-k)/3,1/2\}$, and $w=T^{q}$ with $q=\min\{2k/3,1\}$. In addition, our results provide tighter regret bounds for many applications where system designers possess some knowledge of the transition kernel structure beforehand and can be effectively to problems exhibiting stationary and non-stationary dynamics.
  • Our simulation results show that our algorithm achieves sub-linear regret in the total number of episodes. They also show that our algorithm achieves lower regret in practice compared with the UCWhittle policy [23], WIQL policy [27], and a uniformly randomized policy [3] baselines.
2 Related Work
Offline Whittle Index Policy for RMABs: Whittle’s seminal work [1] introduced a heuristic policy for the infinite-horizon RMAB problem, known as the Whittle index policy. This policy relies on establishing a special mathematical property called indexability for each arm and then deriving functions called index functions that map states to how valuable it would be to activate an arm at that state. Running the policy simply requires activating the $M$ bandits with the highest Whittle indices out of the $N$ bandits at each decision time. The Whittle index achieves asymptotic optimality [19, 20] when the indexability condition [1, 28] is satisfied. Motivated by Whittle’s work, many subsequent works have applied the Whittle index framework to different resource allocation problems [21, 4, 5, 22, 12, 3, 6, 29] by modeling them as RMABs.
Online Learning of Whittle Index: When the transition kernels of a RMAB are unknown and stationary, the problem of finding the Whittle index becomes an online/reinforcement learning problem. In this effort, multiple works [30, 31, 27] have proposed Q-learning algorithms to compute Whittle Index. Authors in [32] proposed NeurWIN and [33] proposed DeepTOP to compute Whittle index by using neural networks. These prior works did not provide any regret guarantees for their policy. Evaluating regret with respect to the optimal policy requires computing the optimal solution to the RMAB problem, which is intractable. In [34] the authors develop an online Whittle algorithm with static regret guarantees compared to the best fixed Whittle index policy. However, they only consider time-varying rewards, the transition kernels remain static. Another approach to make the objective tractable is proposed in [23], which defines regret in terms of the relaxed Lagrangian of the RMAB. In this Lagrangian setting, the authors of [23] present UCWhittle, a Whittle index-based algorithm that uses an upper confidence bound approach to learn the Whittle index. The UCWhittle algorithm achieves $O(\sqrt{T\mathrm{log}T})$ for RMAB with unknown and stationary transition kernels. Unlike prior works, we propose a Whittle index-based algorithm that uses a sliding window and an upper confidence bound approach to learn Whittle index for non-stationary transition dynamics, with provable regret bounds. To the best of our knowledge, this is the first work to provide dynamic regret analysis of an online Whittle index-based policy for RMABs with unknown and non-stationary transition dynamics.
3 Problem Setting
We consider an episodic RMAB problem with $N$ arms and an unknown non-stationary environment. Each arm $n\in[N]$ is associated with a unichain MDP denoted by a tuple $(\mathcal{S},\mathcal{A},P_{n,t},r_{n})$ at every episode $t$, where the state space $\mathcal{S}$ is finite, $\mathcal{A}=\{0,1\}$ is a set of binary actions, $P_{n,t}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ is the transition kernel of arm $n$ with $P_{n,t}(s^{\prime}_{n}|s_{n},a_{n})$ being the probability of transitioning to state $s^{\prime}_{n}$ from state $s_{n}$ by taking action $a_{n}$ in episode $t$, and $r_{n}(s_{n},a_{n})$ is the reward function for arm $n$ when the current state is $s_{n}$ and the action $a_{n}$ is taken. The total number of episodes is $T$ and each episode itself consists of $H$ time slots. We consider that $P_{n,t}$ is unknown and non-stationary, i.e., $P_{n,t}$ can change over episode $t\in[T]$.
A decision maker (DM) determines what action to apply to each arm at a decision time $h\in[H]$ of an episode $t\in[T]$ under the instantaneous activation constraint that at most $M$ arms can be activated. The action taken by the DM in episode $t$ is described by a deterministic policy $\pi_{t}:\mathcal{S}^{N}\mapsto\mathcal{A}^{N}$ which maps a given state $\mathbf{s}=(s_{1},s_{2},\ldots,s_{M})\in\mathcal{S}^{N}$ to an action $\mathbf{a}=(a_{1},a_{2},\ldots,a_{M})\in\mathcal{A}^{N}$. The corresponding expected discounted reward in episode $t$ is given by
$R_{t}\left(\pi_{t},(P_{n,t})_{n=1}^{N}\right):=\mathbb{E}\left[\sum_{h=1}^{H}\sum_{n=1}^{N}\gamma^{h-1}r_{n}(s_{n,h,t},a_{n,h,t})\bigg{|}\pi_{t},(P_{n,t})_{n=1}^{N}\right],$ (1)
where $s_{n,h,t}\in\mathcal{S}$ is the state of arm $n$ at time $h$ of episode $t$, $a_{n,h,t}\in\mathcal{A}$ is the action taken by the DM for arm $n$ at decision time slot $h$ of episode $t$, and $\gamma$ is the discount factor. The DM aims to maximize the total expected sum reward across all episodes, subject to arm activation constraints, i.e.,
$\max_{\pi_{t}\in\Pi}R_{t}\left(\pi_{t},(P_{n,t})_{n=1}^{N}\right);~{}~{}\mathrm{s.t.}\sum_{n=1}^{N}a_{n,h,t}\leq M,\forall h\in[H],\forall t\in[T]$ (2)
where $\Pi$ is the set of all causal policy $\pi_{t}:\mathcal{S}^{N}\mapsto\{0,1\}^{N}$ that maps the state $\mathbf{s}\in\mathcal{S}^{N}$ to $\mathbf{a}\in\mathcal{A}^{N}$.
3.1 Lagrangian Relaxation
Because the main problem (2) is intractable, we relax the per-time slot constraint and use the Lagrangian defined below:
$\mathbb{E}\left[\sum_{h=1}^{H}\sum_{n=1}^{N}\gamma^{h-1}\bigg{(}r_{n}(s_{n,h,t},a_{n,h,t})-\lambda a_{n,h,t}\bigg{)}\bigg{|}\pi_{t},(P_{n,t})_{n=1}^{N}\right],$ (3)
where $\lambda\geq 0$ is a Lagrangian penalty that is interpreted as the cost to pay for activation.
The Lagrangian problem (3) enables us to decompose the combinatorial decision problem (2) into a set of $N$ independent Markov decision process for each arm:
$U\left(\pi^{*}_{n,t},P_{n,t},\lambda\right)=\max_{\pi_{n,t}\in\Pi_{n}}\mathbb{E}\left[\sum_{h=1}^{H}\gamma^{h-1}\bigg{(}r_{n}(s_{n,h,t},a_{n,h,t})-\lambda a_{n,h,t}\bigg{)}\bigg{|}\pi_{n,t},P_{n,t}\right],$ (4)
where we optimize policy $\pi_{n,t}:\mathcal{S}\mapsto\mathcal{A}$ for each arm $n$ from the set of all causal policies $\pi_{n,t}\in\Pi_{n}$.
3.2 Whittle Index Policy
Given $\lambda$, we denote $\phi_{n}(\lambda)$ by the set of states for which it is optimal to activate the arm. The set $\phi_{n}(\lambda)$ is given by $\phi_{n}(\lambda):=\{s\in\mathcal{S}:Q_{n,\lambda,t}(s,0)>Q_{n,\lambda,t}(s,1)\}$, where the action value function $Q_{n,\lambda,t}(s,a)$ associated with Bellman optimality equation for (4) is
$Q_{n,\lambda,t}(s,a)=r_{n}(s,a)-\lambda a+\gamma\sum_{s^{\prime}\in\mathcal{S}}P_{n,t}(s^{\prime}|s,a)V_{n,\lambda,t}(s^{\prime})$ (5)
and the value function $V_{n,\lambda,t}(s)$ associated with Bellman optimality equation for (4) is
$V_{n,\lambda,t}(s)=\max_{a\in\mathcal{A}}Q_{n,\lambda,t}(s,a).$ (6)
Intuitively, as the Lagrangian cost $\lambda$ increases, it is less likely the optimal policy activates arm $n$ in a given state. Hence, the set $\phi_{n}(\lambda)$ would increase monotonically.
Definition 1 (Indexability).
An arm is said to be indexable if the set $\phi_{n}(\lambda)$ increases monotonically as $\lambda$ increases from 00 to $\infty$. A restless bandit problem is said to be indexable if all arms are indexable.
Definition 2 (Whittle Index).
Given indexablity and transition kernel $P_{n,t}$, the Whittle index $W_{n,t}(s;P_{n,t})$ of arm $n$ at state $s\in\mathcal{S}$ in episode $t$ is defined as:
$W_{n,t}(s;P_{n,t}):=\inf\{\lambda:Q_{n,\lambda,t}(s,0)=Q_{n,\lambda,t}(s,1)\}.$ (7)
The Whittle index $W_{n,t}(s;P_{n,t})$ represents the minimum activation cost at which activating arm $n$ in state $s$ at episode $t$ is equally optimal to not activating it.
Whittle Index Policy activates at most $M$ arms out of $N$ arms with highest Whittle indices. However, as we can observe from (7), we can compute Whittle index if we know the transition kernel $P_{n,t}$ of every episode $t\in[T]$. Next, we model how transition kernels change over every episode.
3.3 Transition Kernels Model
In this section, we model the non-stationarity of transition kernels in RMABs. Since transition kernels in an RMAB may drift at varying rates across different arms $n\in[N]$, we define an arm-specific upper bound, $\epsilon_{n}$, on the rate at which the transition kernel of arm $n$ can change over two consecutive episodes.
We consider that the total variation distance [35] between transition kernels in two consecutive episodes is upper bounded by a parameter $\epsilon_{n}$ for arm $n$. Specifically, we have
$\max_{(s,a)\in\mathcal{S}\times A}\sum_{s^{\prime}\in\mathcal{S}}\|P_{n,t}(s^{\prime}|s,a)-P_{n,t-1}(s^{\prime}|s,a)\|_{1}\leq\epsilon_{n}.$ (8)
The parameter $\epsilon_{n}$ determines how quickly the transition kernels change. If $\epsilon_{n}\rightarrow 0$, then transition kernels of arm $n$ are close to stationary, whereas if $\epsilon_{n}$ is far from 00, the transition kernels of arm $n$ become highly non-stationary.
In an RMAB problem, stationary and non-stationary transition kernels can co-exist. To address this co-existing dynamics, we categorize the state-action pairs for each arm $n\in[N]$ into the following two sets based on prior domain knowledge:
$\mathcal{Z}_{1}^{(n)}$ =$=\{(s,a)\in\mathcal{S}\times\mathcal{A}:P_{n,t}(s^{\prime}|s,a)=q(s^{\prime},a,s)~{}\text{is stationary but unknown},~{}\forall s^{\prime}\in\mathcal{S}\},$ (9)
$\mathcal{Z}_{2}^{(n)}$ =$=\{(s,a)\in\mathcal{S}\times\mathcal{A}:P_{n,t}(s^{\prime}|s,a)~{}\text{is non-stationary and unknown},~{}\forall s^{\prime}\in\mathcal{S}\}.$ (10)
If $|\mathcal{Z}^{(n)}_{1}|=0$ and $|\mathcal{Z}^{(n)}_{2}|=|\mathcal{S}\times\mathcal{A}|$, then the transition probabilities for all state-action pairs in arm $n$ are non-stationary. On the other hand, If $|\mathcal{Z}^{(n)}_{1}|=|\mathcal{S}\times\mathcal{A}|$ and $|\mathcal{Z}^{(n)}_{2}|=0$, then the transition probabilities for all state-action pairs in arm $n$ are stationary.
In many applications, the probability transition kernels are sparse - meaning that many state transitions are not possible under certain actions. To model this we introduce $\mathcal{S}_{0}^{(n)}(s,a)$ as the set of all states $s^{\prime}\in\mathcal{S}$ such that the probability to transit from state $s\in\mathcal{S}$ to state $s^{\prime}\in\mathcal{S}$ given action $a\in\mathcal{A}$ is always 00, i.e.,
$\mathcal{S}_{0}^{(n)}(s,a)$ =$=\{s^{\prime}\in\mathcal{S}:P_{n,t}(s^{\prime}|s,a)=0,\forall t\}.$ (11)
The sets $\mathcal{S}_{0}^{(n)}(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$ represents the sparsity of transition kernels for arm $n$.
3.4 Problem Statement
The DM is assumed to know the parameter $\epsilon_{n}$, the sets $\mathcal{Z}^{(n)}_{1}$, $\mathcal{Z}^{(n)}_{2}$, and $\mathcal{S}_{0}^{(n)}(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. The goal of the DM is to design an online algorithm that runs over a sequence of episodes and maximizes the sum of discounted rewards across all episodes. Towards this end, we develop our online algorithm that learns the probability transition kernels and uses the Whittle Index to pick approximately optimal policies in each episode. In the next section, we provide our online algorithm.
4 Sliding Window-based Online Whittle Policy
Recall from earlier discussion (7) that to compute the Whittle index, we need to know transition kernels. In practice, transition kernels $P_{n,t}$ are unknown and non-stationary. In this section, we present Algorithm 1, an online approach for RMABs which adapts to unknown and non-stationary transition kernels.
input : State Space $\mathcal{S}$, Action Space $\mathcal{A}$, Reward Function $r_{n}(s,a)$, and the other domain knowledge $\epsilon_{n}$, $Z_{1}^{(n)},Z_{2}^{(n)},$ and $S_{0}^{(n)}(s,a)$ for all $(s,a)$ and arms $n$
1 DM decides window size $w_{n}$ for all $n\in[N]$, $\eta_{1}$, $\eta_{2}$ and initialize a Lagrange cost $\lambda^{(1)}$
2 for every episode $t=1,2,\ldots,T$ do
3      All arms starts with state $s_{n,0}$ for all $n\in[N]$
4       DM predicts $\tilde{P}_{n,t}$ for all arm $n\in[N]$ using (15) with $\lambda^{(t)}$.
5       DM computes Whittle Index $W_{n,t}(s)\forall s\in\mathcal{S},n\in[N]$ with $\tilde{P}_{n,t}$ using (7).
6       for $h=1,2,\ldots,H$ do
7            DM activates $M$ arms (i.e., action=1) with highest Whittle Indices $W_{n,t}(s_{n,t,h})$.
8             All arms $n$ moves to the next state $s_{n,h+1,t}\sim P_{n,t}(\cdot|s_{n,h,t},a_{n,h,t})$
9             DM observes states $\{s_{n,h+1,t},\forall n\in[N]\}$
10             DM updates counts $C_{t,w_{n}}^{(n)}(s_{n,h+1,t},s_{n,h,t},a_{n,h,t})$ if $(s_{n,h,t},a_{n,h,t})\in\mathcal{Z}_{2}^{(n)}$, otherwise if $(s_{n,h,t},a_{n,h,t})\in\mathcal{Z}_{1}^{(n)}$, DM updates $C_{t,t-1}^{(n)}(s_{n,h+1,t},s_{n,h,t},a_{n,h,t})$
11      Update $\lambda^{(t+1)}=$ M-th highest Whittle Index
Algorithm 1 Sliding Window-based Online Whittle Policy
Our sliding window-based online Whittle policy, provided in Algorithm 1, is motivated by the UCWhittle approach proposed in [23]. However, the UCWhittle policy is designed for static settings and does not handle time-varying transition kernels. This motivates the two main technical innovations in our policy. First, we employ a sliding window method that tracks transition kernels of the past $w$ episodes instead of all past episodes (with different sliding windows for different parts of the state space). Second, we utilize two distinct confidence bounds for stationary sets $\mathcal{Z}^{(n)}_{1}$ and non-stationary sets $\mathcal{Z}^{(n)}_{2}$. The window size and the confidence bound for non-stationary sets $\mathcal{Z}^{(n)}_{2}$ are designed as a function of the parameter $\epsilon_{n}$, the upper bound of the divergence rate of transition kernels.
4.1 Sliding Window and Confidence Bounds
At each episode $t$ and for each arm $n$, we maintain variables $C_{t,w}^{(n)}(s^{\prime},a,s)$, which count the number of transitions from state $s$ to the state $s^{\prime}$ via the action $a$ observed within the past $w$ episodes, i.e. the sliding window. By using the counts for past $w$ episodes, we compute the empirical transition probabilities
$\hat{P}_{n,t,w}(s^{\prime}|s,a):=\frac{C_{t,w}^{(n)}(s^{\prime},a,s)}{C^{(n)}_{t,w}(s,a)},$ (12)
where we define $C^{(n)}_{t,w}(s,a):=\max\bigg{\{}\sum_{s^{\prime}\in\mathcal{S}}C_{t,w}^{(n)}(s^{\prime},a,s),1\bigg{\}}.$
Unlike other upper confidence bound approach [23], we consider two different types of confidence radius; the first one is for the set of states which have unknown but stationary transition probabilities ($\mathcal{Z}_{1}^{(n)}$) and the second one is for states with non-stationary transition probabilities ($\mathcal{Z}_{2}^{(n)}$). We represent these confidence radii as follows
$d_{t,1}^{n}(s,a)=\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}_{1}^{(n)}|NT/\eta_{1})}{C^{(n)}_{t,t-1}(s,a)}},\quad d_{t,2}^{n}(s,a)=\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}_{2}^{(n)}|NT/\eta_{1})}{C^{(n)}_{t,w}(s,a)}}+w\epsilon_{n},$ (13)
where $\eta_{1},\eta_{2}>0$, and the sliding window $w\geq 1$ are design parameters. Notice that the term $w\epsilon_{n}$ in the confidence radius $d_{t,2}^{n}(s,a)$ for $\mathcal{Z}^{(n)}_{2}$ measures how far the transition kernels could have drifted over a window of $w$ episodes.
Equipped with these definitions, the ball $B_{t}^{(n)}$ of the possible values for transition probabilities $P_{n,t}(s^{\prime}|s,a)$ at any episode $t$ can be characterized as follows
$B_{t}^{(n)}=\bigg{\{}P_{n,t}:\sum_{s^{\prime}\in\mathcal{S}}$ $\bigg{|}P_{n,t}(s^{\prime}|s,a)-\hat{P}_{n,t,w_{i}}(s^{\prime}|s,a)\bigg{|}\leq d_{t,i}^{(n)}(s,a),~{}\forall(s,a)\in\mathcal{Z}^{(n)}_{i},i\in\{1,2\},$
$w_{1}=t,w_{2}=w,P_{n,t}(s^{\prime}|s,a)=0,\forall s^{\prime}\in\mathcal{S}^{(n)}_{0}(s,a),\forall(s,a)\in\mathcal{S}\times\mathcal{A}\bigg{\}}.$ (14)
We will show later that the true transition kernel lies within this high-dimensional ball with high probability in each episode.
4.2 Online Whittle Indices
Similar to [23], we predict the transition probabilities in an optimistic approach. We select the optimistic transition probability $\tilde{P}_{n,t}$ for each arm $n$ that maximizes the value function within the confidence bound. The optimization problem for predicting the transition probability $\tilde{P}_{n,t}$ is given by
$\max_{P_{n,t}\in B^{(n)}_{t}}V_{n,\lambda,t}(s),~{}\mathrm{s.t.}$ $~{}V_{n,\lambda,t}(s)=\max_{a\in\mathcal{A}}Q_{n,\lambda,t}(s,a),$ (15)
$Q_{n,\lambda,t}(s,a)=r_{n}(s,a)-\lambda a+\sum_{s^{\prime}}P_{n,t}(s^{\prime}|s,a)V_{n,\lambda,t}(s^{\prime})$ (16)
As a result of the maximization procedure (15), the true value function is upper bounded by the value function under the predicted transition kernel provided that the confidence bound (4.1) holds. This upper bound value function will later alllow us to prove regret bounds. Using the predicted transition kernel $\tilde{P}_{n,t}$, we compute $W_{n,t}(s:\tilde{P}_{n,t})$, the Whittle index of state $s\in\mathcal{S}$ for arm $n$ as defined in (7). Finally, we update Lagrange multiplier $\lambda^{t+1}$ as the $M$-th highest Whittle index at time slot $H$ of episode $t$. However, the approach can be computationally intensive due to the kernel maximization problem (15). Some approaches to reduce the complexity are discussed in Appendix D.
5 Regret Analysis
It is well known in RMAB literature, the problem of obtaining an optimal policy for an RMAB problem is generally intractable [18]. Consequently, using an optimal policy for an RMAB is an impractical benchmark to evaluate the online learning performance. To address this issue, [23] and [34] used the Lagrangian relaxed problem or alternatively the best Whittle index policy to assess the performance of an online learning algorithm. Following [23], we will also employ the Lagrangian relaxed problem for evaluating the online learning performance in this paper.
We determine the regret of the policy $\pi_{t}$ in episode $t$ by subtracting the performance of our policy from the performance of the optimal policy (both under the true unknown transition kernel $P_{n,t}$) for the Lagrange relaxed problem.
$\mathrm{Reg}^{(t)}=\sum_{n=1}^{N}U(\pi^{*}_{n,t},P_{n,t},\lambda^{*})-U(\pi_{n,t},P_{n,t},\lambda^{(t)}),$ (17)
where $\pi^{*}_{n,t}$ is the optimal policy of the Lagrange relaxed problem (4) and $\lambda^{*}$ minimizes Lagrangian objective value $\sum_{n=1}^{N}U\left(\pi^{*}_{n,t},P_{n,t},\lambda\right)$. The cumulative dynamic regret in $T$ episodes is given by $\mathrm{Reg}(T)=\sum_{t=1}^{T}\mathrm{Reg}^{(t)}.$ Note that sublinear dynamic regret is usually challenging to establish in online learning literature, since we are comparing to a dynamic optimal policy that knows the entire sequence of transition kernels [36, 37]. We will show that our approach has sublinear dynamic regret, as long as the transition kernels don’t vary too quickly.
To create a regret bound, we first need to establish how good our estimates of the time-varying transition kernel are. To do so, we will bound the probability that the true kernel is outside the high-dimensional ball $B_{t}^{(n)}$ introduced in (4.1). Lemma 1 describes the result in detail.
Lemma 1.
Given $\epsilon_{n},\eta_{1},\eta_{2}\geq 0$, the probability that the true kernel $P_{n,t}$ lies within this high-dimensional Ball $B_{t}^{(n)}$ is greater or equal to $1-\eta$, where $\eta=\eta_{1}+\eta_{2}$, i.e., $\mathrm{Pr}(P_{n,t}\in B_{n,t},\forall n,\forall t)\geq 1-\eta$.
Lemma 1 implies that for every episode $t$, we can provide a confidence region in which true transition kernel will lie with high probability. A detailed proof of Lemma 1 is provided in Appendix A. Next, using this result, Theorem 1 characterizes the dynamic regret for our proposed Algorithm 1.
Theorem 1.
With probability $1-\eta$, the cumulative dynamic regret of Algorithm 1 satisfies:
$\mathrm{Reg}(T)\leq O\left(\bigg{(}\sum_{n=1}^{N}|\mathcal{Z}^{(n)}_{1}|\bigg{)}H\sqrt{T\mathrm{log}T}\right)+\sum_{t=1}^{T}O\bigg{(}\sum_{n=1}^{N}|\mathcal{Z}^{(n)}_{2}|\bigg{(}G_{t,n}(w_{n})+w_{n}\epsilon_{n}\bigg{)}\bigg{)},$ (18)
where $G_{t,n}(w)=\max_{(s,a)\in\mathcal{Z}_{2}}g_{t,n}(s,a,w)$ and the function $g_{t,n}(s,a,w)$ is a non-decreasing function given by $g_{t,n}(s,a,w)=\mathbb{E}_{P_{n,t},\pi_{n,t}}[\alpha_{t}^{(n)}(s,a)/\sqrt{C_{t,w}^{(n)}(s,a)}]$, where $\alpha_{t}^{(n)}(s,a)$ is a random variable that denotes the number of visit at $(s,a)$ in episode $t$.
Proof Sketch. We focus on bounding the dynamic regret when the confidence bound holds.
$\mathrm{Reg}^{(t)}$ $=U(\pi^{*}_{n,t},P_{n,t},\lambda)-U(\pi_{n,t},P_{n,t},\lambda)\overset{(a)}{\leq}U(\pi_{n,t},\tilde{P}_{n,t},\lambda)-U(\pi_{n,t},P_{n,t},\lambda)$
$\overset{(b)}{\leq}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{h=1}^{H}\sum_{(s,a)\in\mathcal{S}}\alpha_{t}^{(n)}(s,a)\sum_{s^{\prime}\in\mathcal{S}}\bigg{|}\tilde{P}_{n,t}(s^{\prime}|s,a)-P_{n,t}(s^{\prime}|s,a)\bigg{|}\bigg{]}V_{max},$ (19)
where (a) holds due to (15) and (b) is obtained by using the random variable $\alpha_{t}^{(n)}(s,a)$. Then, we utilize the confidence radii $d_{t,1}^{(n)}(s,a)$, $d_{t,2}^{(n)}(s,a)$, and sum the regret $\mathrm{Reg}^{(t)}$ over episodes $t=1$ to $T$ to get the final regret bound. A detailed proof is provided in Appendix B.
At a high-level, the regret bound consists of two terms - one for parts of the transition kernels that are stationary, and one for parts that are non-stationary. The second term grows with $\epsilon_{n}$ and depends crucially on the choice of window sizes for each arm. Moreover, as the size of the sets $\mathcal{Z}^{(n)}_{1}$ and $\mathcal{Z}^{(n)}_{2}$ increases, the regret increases. If the transition kernels are stationary, i.e., $\epsilon=0$, then our regret becomes $O((\sum_{n=1}^{N}|\mathcal{Z}^{(n)}_{1}|)H\sqrt{T\mathrm{log}T}),$ which is better compared to UCWhittle’s [23] regret upper bound $O\left(NH|\mathcal{S}|\sqrt{T\mathrm{log}T}\right)$ since $\mathcal{Z}^{n}_{1}\subseteq\mathcal{S}$. This shows that prior domain knowledge is helpful both for stationary and non-stationary settings.
While Theorem 1 characterizes the regret of our algorithm, the following theorem describes how to optimally choose window sizes for each arm. We are able to show that for slowly varying kernels, our algorithm has sublinear dynamic regret.
Theorem 2.
If there exists a positive probability to visit every $(s,a)\in\mathcal{Z}_{2}$ at least once in every episode $t\in[T]$ for all arms $n\in[N]$ and the divergence of transition dynamics between two consecutive episodes is upper bounded by $\epsilon_{n}=1/T^{k}$, then by using window size $w_{n}=T^{q}$ and $q=\min\{2k/3,1\}$, with probability $1-\eta$, we have $\mathrm{Reg}(T)\leq O(H\sqrt{\mathrm{log}(T)}T^{r})$, where $r=\max\{(3-k)/3,1/2\}$.
Proof Sketch. Given our assumptions, we can upper bound the function $G_{t,n}(w_{n})$ provided in Theorem 1 by $\frac{H}{\sqrt{\mathrm{P}_{min}w}}$, where $\mathrm{P}_{min}$ is the minimum probability to visit any state-action $(s,a)\in\mathcal{Z}_{2}$ at least once at every episode $t\in[T]$. Inserting this bound in Theorem 1 yields Theorem 2. A detailed proof is provided in Appendix C.
6 Numerical Study
Applications $N$ $M$ Our Policy UCWhittle Random WIQL
[…]
10 4 2713±$\pm 261$ 61404 ±$\pm 175$ 83668 ±$\pm 479$ 65694±$\pm 2305$
20 4 1954±$\pm 198$ 107610 ±$\pm 176$ 180770 ±$\pm 321$ 18110-±$\pm 2122$
[…]
10 4 426.46±$\pm 23$ 958.96±$\pm 63$ 4579.8±$\pm 39$ 5172.2±$\pm 37$
20 4 656.53±$\pm 41$ 2637.1±$\pm 155$ 13609±$\pm 178$ 12584±$\pm 137$
Table 1: $\mathrm{Reg}(T)$ for $T=50$ episodes for different values of $N$ and $M$ with $H=100$ and $\epsilon_{n}=0.05$
In this section, we demonstrate the performance of our proposed policy by evaluating it under two applications modeled as restless multi-armed bandit problems. In each application, we consider that there are $N$ arms and a policy can activate $M$ of them in each time slot $h\in[H]$ of every episode $t\in[T]$. We evaluate our policy against the UCWhittle policy [23], the WIQL policy [27], and a uniformly randomized policy [3]. The resultsare averaged over $50$ independent runs.
One Dimensional Bandits: We consider a modified version of the one dimensional RMAB problem studied in [38, 33]. Each arm $n$ is a Markov process with $K$ states, numbered as $10,1,\ldots,K-$ 1. For our simulations, we set $K=10$. The reward of an arm increases linearly with the current state, i.e. $r(s,a)=s$. If the arm is activated, then it can evolve from state $s$ to $\min\{s+1,K-1\}$ with probability $q_{n}(t)$ or remain in the same state $s$ with probability $1-q_{n}(t)$. If the arm is not activated, then it evolves from state $s$ to $\max\{s-1,0\}$ with probability $p_{n}(t)$ or remain in the same state $s$ with probability $1-p_{n}(t)$. One-dimensional MDPs of this form are often used in health monitoring and machine monitoring applications [39, 40]. In our simulation, for half of the $N$ arms, we consider the following kernel dynamics which are unknown and must be learned from data - (i) $p_{n}(t)$ changes to $\min\{p_{n}(t-1)+\epsilon_{n},1\}$ with probability $0.7$, or it changes to $\max\{p_{n}(t-1)-\epsilon_{n},0\}$ with probability 0.3 and (ii) $q_{n}(t)$ is unknown but stationary. For the other half, probabilities $p_{n}(t)$ and $q_{n}(t)$ are known to the decision maker and do not change over episode $t$. The detailed experimental setup is provided in the Appendix D.1.
Wireless Scheduling: We consider a wireless scheduling problem, where $M$ out of $N$ sources can send their observation to a receiver side over an unreliable channel at every time slot $h\in[H]$ of episode $t\in[T]$. Due to channel unreliability, the observation may not be delivered. The goal of the receiver is to estimate the current signal values of all $N$ sources based the information delivered from the sources. The reward for accurate timely estimation can be modeled as the mutual information between the estimated signal and the actual signal. [41] showed that the mutual information can be determined by using a decreasing function $-(\mathrm{log}_{2}(1-\sigma_{n}^{2s_{n,h,t}}))/2$ of Age of Information (AoI) for zero-mean i.i.d. Gaussian random variables with variance $\sigma^{2}_{n}$, where AoI $s_{n,h,t}$ of source $n$ is the time difference between current time $h$ and the generation time of the most recently delivered signal. The AoI value of a source $n$ increases by $1$ if the source $n$ is not scheduled. If the source $n$ is scheduled, the AoI value drops to $1$ with probability $q_{n}(t)$ (successful delivery) or increases by $1$ with probability $1-q_{n}(t)$ (unsucessful delivery). The parameter $q_{n}(t)$ measures the reliability of channel $n$ at time $t$. In our experiment, we assume that $q_{n}(t)$ is unknown and non-stationary for half of the sources, whereas it is is unknown but stationary for the remaining half. The detailed experimental setup is provided in the Appendix D.2.
Figure 1: One Dimensional Bandit: Cumulative regret versus the number of episodes with different values of time horizon ($H=20,50,100$) with $\epsilon_{n}=0.05$.
Figure 2: Wireless Scheduling: Cumulative regret versus the number of episodes with different values of time horizon ($H=20,50,200$) with $\epsilon_{n}=0.05$.
Simulation results: Simulation results are shown in Table 1, Figure 1 and Figure 2. It can be observed that our algorithm achieves the best regret in all cases. UCWhittle [23] is designed for static settings, hence it uses all historical data to learn transitions, which often yields inaccurate transition kernels for the current episode. Moreover, UCWhittle naively attempts to learn the entire transition matrix. The stationary random policy [3] picks $M$ out of $N$ arms uniformly at random, clearly leading to worse performance. WIQL [27], a Q-learning approach, requires many data samples to converge. Also, similar to UCWhittle, WIQL uses all historical data. Due to WIQL’s slow convergence, WIQL shows similar performance compared to random policy, see [23]. Because of the non-stationary setting, WIQL performs poorly even compared to random policy, suggesting that traditional reinforcement learning methods cannot be naively applied to solve dynamic RMAB problems.
7 Conclusions and Limitations
This paper introduced an online/reinforcement learning algorithm for estimating the Whittle index for restless bandit problems with unknown and non-stationary transition kernels using sliding window and upper confidence bound approaches. To our knowledge, this is the first work to provide an upper bound of the dynamic regret of an online Whittle index-based algorithm for RMABs with unknown and non-stationary transition kernels. Our proposed algorithm is evaluated on two different restless bandit problems against three baselines and provides significant performance gains. We also provide novel regret analysis and tighter regret bounds for RMABs in both stationary and non-stationary settings. An interesting direction of future work involves proving lower bounds for regret, i.e., showing that no online algorithm can achieve regret better than $O(\sqrt{T\mathrm{log}(T)})$. Moreover, our algorithm can only be applied to a finite state space and works best for small $\epsilon_{n}$ (slowly varying kernels). Promising future directions include extending this work to infinite or continuous state spaces, and designing algorithms that achieve sub-linear dynamic regret even for large $\epsilon_{n}$ (rapidly varying kernels).
Appendix A Proof of Lemma 1
We can get Lemma 1 by combining Lemma 2 and Lemma 3, which are provided below:
Lemma 2.
Given $\eta_{1}>0$ and $t\geq 1$, for $(s,a)\in\mathcal{Z}^{(n)}_{1}$, we have
$\mathrm{Pr}\left(\|\hat{P}_{n,t,t-1}(\cdot|s,a)-P_{n,t}(\cdot|s,a)\|_{1}\geq\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}_{1}^{(n)}|NT/\eta_{1})}{C^{(n)}_{t,t-1}(s,a)}}\right)\leq\frac{\eta_{1}}{N|\mathcal{Z}^{(n)}_{1}|T}.$ (20)
Proof.
The L1-deviation of the true distribution and the empirical distribution of $m$ events is bounded by [42]:
$\mathrm{Pr}(|\hat{p}-p|_{1}\geq\beta)\leq(2^{m}-2)\exp^{(-\frac{n\beta^{2}}{2})},$ (21)
where $n$ is the number of samples.
By applying (21) with
$\beta=\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}_{1}^{(n)}|NT/\eta_{1})}{C^{(n)}_{t,t-1}(s,a)}}$
and
$n=C^{(n)}_{t,t-1}(s,a),$
we have
$\mathrm{Pr}\left(\|\hat{P}_{n,t,t-1}(\cdot|s,a)-P_{n,t}(\cdot|s,a)\|_{1}\geq\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}_{1}^{(n)}|NT/\eta_{1})}{C^{(n)}_{t,t-1}(s,a)}}\right)$
$\leq 2^{|\mathcal{S}|}\exp(-\mathrm{log}(2^{|\mathcal{S}|}|\mathcal{Z}_{1}^{(n)}|NT/\eta_{1}))=\frac{\eta_{1}}{N|\mathcal{Z}_{1}^{(n)}|T}.$ (22)
Lemma 3.
Given $\eta_{2}>0$ and $w\geq 1$, for $(s,a)\in\mathcal{Z}_{2}$, we have
$\mathrm{Pr}\left(\|\hat{P}_{n,t,w}(\cdot|s,a)-P_{n,t}(\cdot|s,a)\|_{1}\geq\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}_{2}^{(n)}|NT/\eta_{2})}{C^{(n)}_{t,w}(s,a)}}+w\epsilon\right)\leq\frac{\eta_{2}}{N|\mathcal{Z}^{(n)}_{2}|T}.$ (23)
Proof.
We denote $\mathbf{1}(s^{\prime},s,a,n,t)$ as an indicator variable that represents the event of state $s$, action $a$, and next state $s^{\prime}$ for arm $n$ at one time slot of episode $t$. Similarly, $\mathbf{1}(s^{\prime},s,a,n,t,w)$ is an indicator variable that represents the event of state $s$, action $a$, and next state $s^{\prime}$ for arm $n$ at one time slot in any one of the episodes $t-w+1,t-(w-1)+1,\ldots,t-1$.
By using (21) with
$\beta=\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}^{(n)}_{2}|NT/\eta_{2})}{C^{(n)}_{t,w}(s,a)}}$
and
$n=C^{(n)}_{t,w}(s,a),$
we have
$\mathrm{Pr}\left(\|\hat{P}_{n,t,w}(\cdot|s,a)-\mathbb{E}[\mathbf{1}(\cdot,s,a,n,t,w)]\|_{1}\geq\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}^{(n)}_{2}|NT/\eta_{2})}{C^{(n)}_{t,w}(s,a)}}\right)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\leq\frac{\eta_{2}}{N|\mathcal{Z}^{(n)}_{2}|T}.$ (24)
With probability one, we have
$\|P_{n,t}(\cdot|s,a)-\mathbb{E}[\mathbf{1}(\cdot,s,a,n,t,w)]\|_{1}$
$\leq$ $\|P_{n,t}(\cdot|s,a)-\max_{t^{\prime}\in\{t-w+1,t-(w-1)+1,\ldots,t\}}\mathbb{E}[\mathbf{1}(\cdot,s,a,n,t^{\prime},1)]\|_{1}$
$=$ $\|P_{n,t}(\cdot|s,a)-\max_{t^{\prime}\in\{t-w+1,t-(w-1)+1,\ldots,t\}}P_{n,t^{\prime}}(\cdot|s,a)]\|_{1}\leq w\epsilon.$ (25)
By combining (A) and (A), we obtain Lemma 3. ∎
Now, by combining Lemma 2 and Lemma 3, we have
$\mathrm{Pr}(P_{n,t}\in B_{n,t},\forall n,\forall t)$ $\geq 1-\sum_{t=1}^{T}\sum_{n=1}^{N}\sum_{(s,a)\in\mathcal{Z}^{(n)}_{1}}\frac{\eta_{1}}{N|\mathcal{Z}_{1}^{(n)}|T}-\sum_{t=1}^{T}\sum_{n=1}^{N}\sum_{(s,a)\in\mathcal{Z}^{(n)}_{2}}\frac{\eta_{2}}{N|\mathcal{Z}_{2}^{(n)}|T}$
$=1-\eta_{1}-\eta_{2}.$ (26)
This concludes the proof of Lemma 1.
Appendix B Proof of Theorem 1
When the confidence bound holds, we have
$\mathrm{Reg}^{(t)}$ $=\sum_{n=1}^{N}U(\pi^{*}_{n,t},P_{n,t},\lambda)-U(\pi_{n,t},P_{n,t},\lambda){\leq}\sum_{n=1}^{N}U(\pi_{n,t},\tilde{P}_{n,t},\lambda)-U(\pi_{n,t},P_{n,t},\lambda)$
$\leq\sum_{n=1}^{N}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}}\gamma^{h-1}(\tilde{P}_{n,t}(s^{\prime}|s_{n,t,h},a_{n,t,h})-P_{n,t}(s^{\prime}|s_{n,t,h},a_{n,t,h}))V_{n}(s^{\prime};\pi_{n,t},P_{n,t})\bigg{]},$
$\leq\sum_{n=1}^{N}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{(s,a)\in\mathcal{S}}\alpha_{t}^{(n)}(s,a)\sum_{s^{\prime}\in\mathcal{S}}\bigg{|}\tilde{P}_{n,t}(s^{\prime}|s,a)-P_{n,t}(s^{\prime}|s,a)\bigg{|}\bigg{]}V_{max},$
$=\sum_{n=1}^{N}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{(s,a)\in\mathcal{Z}^{n}_{1}}\alpha_{t}^{(n)}(s,a)d_{t,1}^{(n)}(s,a)+\sum_{(s,a)\in\mathcal{Z}^{(n)}_{2}}\alpha_{t}^{(n)}(s,a)d_{t,2}^{(n)}(s,a)\bigg{]}V_{max},$ (27)
where the simplified notation $V_{n}$ is used instead of $V_{n,\lambda,t}$, $V_{max}=\max_{n\in[N],s\in\mathcal{S}}V_{n}(s^{\prime};\pi_{n,t},P_{n,t})$ and $\alpha_{t}^{(n)}(s,a)$ is a random variable denoting the number of visits of $(s,a)\in\mathcal{S}\times\mathcal{A}$ at episode $t\in[T]$.
By substituting the value of $d_{t,1}^{(n)}(s,a)$, we have
$\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{(s,a)\in\mathcal{Z}_{1}^{(n)}}\alpha_{t}^{(n)}(s,a)d_{t,1}^{(n)}(s,a)\bigg{]}$
$\leq\sqrt{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}_{1}^{(n)}|NT/\eta_{1})}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{(s,a)\in\mathcal{Z}_{1}}\frac{\alpha_{t}^{(n)}(s,a)}{\sqrt{C^{n}_{t,t-1}(s,a)}}\bigg{]}.$ (28)
Moreover, by applying [23, Lemma E.3], we get
$\sum_{t=1}^{T}\sum_{(s,a)\in\mathcal{Z}_{1}}\frac{\alpha_{t}^{(n)}(s,a)}{\sqrt{C^{(n)}_{t,t-1}(s,a)}}\leq(\sqrt{H+1}+1)\sqrt{|\mathcal{Z}_{1}^{(n)}|TH}$ (29)
Hence,
$\sum_{t=1}^{T}\sum_{n=1}^{N}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{(s,a)\in\mathcal{Z}_{1}^{(n)}}\alpha_{t}^{(n)}(s,a)d_{t,1}^{(n)}(s,a)\bigg{]}V_{\max}\leq O\left(\bigg{(}\sum_{n=1}^{N}|\mathcal{Z}^{(n)}_{1}|\bigg{)}H\sqrt{T\mathrm{log}T}\right).$ (30)
On the other hand, by substituting the value of $d_{t,2}^{(n)}(s,a)$, we have
$\sum_{t=1}^{T}\sum_{n=1}^{N}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{(s,a)\in\mathcal{Z}_{2}^{(n)}}\alpha_{t}^{(n)}(s,a)d_{t,2}^{(n)}(s,a)\bigg{]}$
$=\sum_{t=1}^{T}\bigg{(}\sum_{n=1}^{N}\sqrt{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}^{(n)}_{2}|NT/\eta_{1})}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{(s,a)\in\mathcal{Z}_{2}}\frac{\alpha_{t}^{(n)}(s,a)}{\sqrt{C^{(n)}_{t,w_{n}}(s,a)}}\bigg{]}+w_{n}\epsilon_{n}|\mathcal{Z}^{(n)}_{2}|\bigg{)}$
$=\sum_{t=1}^{T}\bigg{(}\sum_{n=1}^{N}\sqrt{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}^{(n)}_{2}|NT/\eta_{1})}\sum_{(s,a)\in\mathcal{Z}^{(n)}_{2}}g_{t,n}(s,a,w_{n})+w_{n}\epsilon_{n}|\mathcal{Z}_{2}|\bigg{)}$
$\leq\sum_{t=1}^{T}\bigg{(}\sum_{n=1}^{N}\sqrt{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}^{(n)}_{2}|NT/\eta_{1})}|\mathcal{Z}^{(n)}_{2}|G_{t,n}(w_{n})+w_{n}\epsilon_{n}|\mathcal{Z}^{(n)}_{2}|\bigg{)}$
$=\sum_{t=1}^{T}O\bigg{(}\sum_{n=1}^{N}|\mathcal{Z}^{(n)}_{2}|\bigg{(}G_{t,n}(w_{n})+w_{n}\epsilon_{n}\bigg{)}\bigg{)},$ (31)
where $G_{t,n}(w)=\max_{(s,a)\in\mathcal{Z}^{(n)}_{2}}g_{t,n}(s,a,w)$ and
$g_{t,n}(s,a,w)=\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{(s,a)\in\mathcal{Z}_{2}}\frac{\alpha_{t}^{(n)}(s,a)}{\sqrt{C^{(n)}_{t,w}(s,a)}}\bigg{]}$
is a non-increasing function of the window size $w$. This is because $C^{(n)}_{t,w}(s,a)$ is a non-decreasing function of $w$. This conludes the proof.
Appendix C Proof of Theorem 2
To prove Theorem 2, we first bound
$g_{t,n}(s,a,w)$ $=\mathbb{E}_{P_{n,t},\pi_{n,t}}\left[\frac{\alpha_{t}^{(n)}(s,a)}{\sqrt{C^{(n)}_{t,w}(s,a)}}\right]\overset{(a)}{\leq}\frac{H}{\sqrt{w_{n}\mathrm{p}_{min}}},$ (32)
where $(a)$ holds because there exists a positive probability $\mathrm{p}_{min}$ to visit every $(s,a)\in\mathcal{Z}_{2}$ at least once in every episode $t\in[T]$ for all arms $n\in[N]$, so $C^{(n)}_{t,w}(s,a)\geq w_{n}\mathrm{p}_{min}$ and it always satisfies that $\alpha_{t}^{(n)}(s,a)\geq H$.
Then, we can have
$\sum_{t=1}^{T}\bigg{(}G_{t,n}(w_{n})+w_{n}\epsilon_{n}\bigg{)}$
$\leq$ $\sum_{t=1}^{T}\bigg{(}\frac{H}{\sqrt{w_{n}\mathrm{p_{min}}}}+w_{n}\epsilon_{n}\bigg{)}.$ (33)
After that, by substituting $\epsilon_{n}=T^{-k}$ and $w_{n}=T^{q}$, we get
$\sum_{t=1}^{T}\bigg{(}\frac{H}{\sqrt{w_{n}\mathrm{p_{min}}}}+w_{n}\epsilon_{n}\bigg{)}=HT^{1-q/2}+HT^{q+1-k}.$ (34)
Further, by substituting $q=\min\{2k/3,1\}$, we have
$HT^{1-q/2}+HT^{q+1-k}\leq HT^{r},$ (35)
where $r=\max\{(3-k)/3,1/2\}$. Hence,
$\sum_{t=1}^{T}\bigg{(}G_{t,n}(w_{n})+w_{n}\epsilon_{n}\bigg{)}\leq HT^{r}.$ (36)
By substituting the above upper bound in Theorem 1, we can get Theorem 2.
Appendix D Experimental Setup
Firstly, we discuss how we compute the regret in our numerical studies. Given a policy, we evaluate the regret of the policy at each episode $t$ by subtracting the total discounted sum reward of all arms obtained by the policy from the total discounted sum reward of all arms received by a Whittle index policy with access to the true transition probabilities. In our simulation, we have used MATLAB.
In all of our experiments, a discount factor $\gamma=0.99$ is considered. Moreover, we consider $2$ class of arms. In each class, the arms share same transition kernels and same reward function in each episode, which is a common setup in the RMAB literature [5, 23, 20]
Our policy can be computationally intensive due to the kernel maximization problem (15). In the next section, along with more experimental setups, we discuss some approaches to reduce the complexity for One Dimensional Bandit and Wireless Scheduling simulations.
D.1 One Dimensional Bandit
In One Dimensional Bandit applications, we can show that $Q_{n,\lambda,t}(s,0)$ is a decreasing function of $p_{n}(t)$, and $Q_{n,\lambda,t}(s,1)$ is an increasing function of $q_{n}(t)$. This is because the reward $r(s,a)$ is an increasing function of state $s$, which results in $V_{n,\lambda,t}(s)$ being an increasing function of state $s$. Using this structure, we can get closed-form solution of the maximization problem (15).
Recall that, in the numerical study of One Dimensional Bandit, for half of the $N$ arms, we consider the following kernel dynamics which are unknown and must be learned from data - (i) $p_{n}(t)$ changes to $\min\{p_{n}(t-1)+\epsilon_{n},1\}$ with probability $0.7$, or it changes to $\max\{p_{n}(t-1)-\epsilon_{n},0\}$ with probability 0.3; the initial value of $p_{n}(t)=0.5$ is used with $\epsilon_{n}=0.05$ and (ii) $q_{n}(t)=0.5$ is unknown but stationary. For the other half, probabilities $p_{n}(t)=0.5$ and $q_{n}(t)=1$ are known to the decision maker and do not change over episode $t$.
D.2 Wireless Scheduling
For our wireless scheduling problem, the value function $V_{n,\lambda,t}(s)$ is an increasing function of $q_{n}(t)$ because the reward function $r(s,a)$ is a decreasing function of the AoI state $s$. This structure allows us to obtain a closed-form solution for the maximization problem (15). Furthermore, a closed-form expression for the Whittle index exists [4]. Though the closed-form expression is obtained for average cost problem, we can use the Whittle index expression from [4] when the discount factor $\gamma\rightarrow 1$. In our experiment, as we have used $\gamma=0.99$, the approximation does not affect the performance of our policy. Consequently, we can determine the Whittle index for our problem using a low-complexity method that exploits these closed-form expressions.
Recall that in the numerical study of wireless scheduling, we assume that $q_{n}(t)$ is unknown and non-stationary for half of the sources, whereas it is is unknown but stationary for the remaining half. For non-stationary arms, the variance of signal values $\sigma_{n}^{2}=0.9$ is used and the probability of successful transmission $q_{n}(t)$ changes to $\min\{q_{n}(t-1)+\epsilon_{n},1\}$ with probability $0.7$, or it changes to $\max\{q_{n}(t-1)-\epsilon_{n},0\}$ with probability 0.3; the initial value of $q_{n}(t)=0.1$ is used with $\epsilon_{n}=0.05$. For the other half, $q_{n}(t)=1$ is unknown but stationary.